OpenStudy (anonymous):

Solve for x X^2-4x+3x-2=28

OpenStudy (akashdeepdeb):

You can also write this as x^2 -x -2 - 28 = 0 right? So x^2 - x - 30 = 0 Now split the middle term! The terms used in splitting should multiply to 30 but add/subtract to the co-eff of x. We have 6 and 5 that multiply to 30 and when subtracted give us 1!!! :D So we write this as x2 - 6x + 5x - 30 = 0 x(x-6) + 5(x-6) = 0 Or (x+5)(x-6) = 0 Understood? :D

OpenStudy (akashdeepdeb):

Now x would be? -5 and 6 Getting me? :D

OpenStudy (anonymous):

Well yes and no. The original question was: Point K is between J and L. If JK = x^2-4x, KL=3x -2 and JL=28, write and solve an equation to find the lengths of JK and KL. The one you solved was my "equation"

OpenStudy (anonymous):